OpenStudy (anonymous):

Zoila's father gave her $30 to spend at the souvenir shop. Zoila already had $10 saved from babysitting. Zoila bought 5 magnets for $2.99 each and a mug for $5.99. Zoila had to pay 7% tax on her purchases. How much did Zoila spend

OpenStudy (texaschic101):

5 magnets for 2.99 = 5(2.99) = 14.95 and the mug costs 5.99 for a total of (14.95 + 5.99) = 20.94 (this is without tax) she paid 7% tax... 20.94 + .07(20.94) = 20.94 + 1.47 = 22.41 Zoila spent $22.41 at the souvenir shop.

OpenStudy (anonymous):

correct you don't need help hunny